# Omnipredicting Single-Index Models with Multi-Index Models

### v1

@@ 4.2 Proof of Theorem 5
$n=\Theta\left(\frac{L^{2}R^{2}}{\varepsilon^{2}}\cdot\left(\frac{\beta^{2}L^{2}\bar{L}^{2}R^{4}}{\varepsilon^{2}}\log^{3}\left(\frac{\beta L^{2}R^{2}}{\varepsilon}\right)+\log\left(\frac{1}{\delta}\right)\right)\right),$
@@
$n=\Theta\left(\frac{L^{2}R^{2}}{\varepsilon^{2}}\cdot\left(\frac{\beta L\bar{L}R^{2}}{\varepsilon}+d\log\left(\frac{\beta L^{2}R^{2}}{\varepsilon}\right)+\log\left(\frac{1}{\delta}\right)\right)\right).$
@@
Throughout this proof, let $\alpha:=\frac{\varepsilon}{6L\bar{L}R^{2}}$, and let
@@
$n=\Theta\left(\frac{L^{2}R^{2}}{\varepsilon^{2}}\cdot\left(\log\frac{\bar{L}R}{\delta\varepsilon}+\frac{L^{2}\bar{L}^{2}R^{4}\beta^{2}}{\varepsilon^{2}}\log^{3}\left(\frac{L^{2}\bar{L}R^{3}\beta}{\varepsilon^{2}}\right)\right)\right),$
@@
$n=\Theta\left(\frac{L^{2}R^{2}}{\varepsilon^{2}}\cdot\left(\log\frac{\bar{L}R}{\delta\varepsilon}+\frac{\beta L\bar{L}R^{2}}{\varepsilon}+d\log\frac{\beta L\bar{L}R^{2}}{\varepsilon}\right)\right)$
@@ 6.2 Generalization
Let $\mathcal{D}$ be an arbitrary distribution over ×$\mathbb{R}\times\{0,1\}$. For any $\delta\in(0,1/3)$, with probability at least $1-\delta$ over the random draw of $n\geq 2$ i.i.d. data points $(x_{1},y_{1}),\ldots,(x_{n},y_{n})$ from $\mathcal{D}$, the output predictor $p:\mathbb{R}\to[0,1]$ from running PAV on the uniform distribution over the $n$ data points555 When we define PAV in Section 2.4, we assume that the domain $\mathcal{X}$ consists of integers $1,\ldots,n$, but the algorithm extends to any finite domain $\{x_{1},\ldots,x_{n}\}\subseteq\mathbb{R}$ by mapping the elements to $1,\ldots,n$ while preserving the ordering. PAV then gives us a non-decreasing predictor $p:\{x_{1},\ldots,x_{n}\}\to[0,1]$, which we can then extrapolate to a non-decreasing step function (i.e., piece-wise constant with at most $n+1$ pieces) over the entire domain $\mathbb{R}$. satisfies the following properties:
@@ B.1 Omniprediction loss from anti-Lipschitzness
In an instance of Model 1, following the notation (10), let $\sigma\in\mathcal{S}_{0,\beta}$, and let $\alpha\in(0,\frac{1}{2LR})$. There exists a $\sigma^{\prime}\in\mathcal{S}_{\alpha,\alpha+(1-2\alpha LR)\beta}$ such that for all $\mathbf{w}\in\mathcal{W}$,
$\mathbb{E}_{(\mathbf{x},y)\sim\mathcal{D}}\left[\left|\ell_{\mathsf{m},\sigma^{\prime}}(\mathbf{w}\cdot\mathbf{x},y)-\ell_{\mathsf{m},\sigma}(\mathbf{w}\cdot\mathbf{x},y)\right|\right]\leq\frac{3\alpha\bar{L}LR^{2}}{2}.$
@@
$\leq\frac{\alpha t^{2}}{2}+\left|\int_{0}^{t}\alpha LR\,\textup{d}\tau\right|\leq\frac{\alpha t^{2}}{2}+\alpha LR\left|t\right|.$
Thus by taking expectations and using the moment bounds from Model 1, the claim follows:
$\mathbb{E}_{\mathbf{x}\sim\mathcal{D}_{\mathbf{x}}}\left[\frac{\alpha(\mathbf{w}\cdot\mathbf{x})^{2}}{2}\right]$ $\leq\frac{\alpha R^{2}}{2}\left\lVert\mathbb{E}_{\mathbf{x}\sim\mathcal{D}_{\mathbf{x}}}\left[\mathbf{x}\mathbf{x}^{\top}\right]\right\rVert_{\textup{op}}\leq\frac{\alpha\bar{L}^{2}R^{2}}{2},$
$\mathbb{E}_{\mathbf{x}\sim\mathcal{D}_{\mathbf{x}}}\left[\alpha LR\left|\mathbf{w}\cdot\mathbf{x}\right|\right]$ $\leq\alpha LR\sqrt{\mathbb{E}_{\mathbf{x}\sim\mathcal{D}}\left[(\mathbf{w}\cdot\mathbf{x})^{2}\right]}\leq\alpha\bar{L}LR^{2}.$
@@
the conclusion follows from Lemma 24, which gives for our range of $\alpha$,
$\mathbb{E}_{(\mathbf{x},y)\sim\mathcal{D}}\left[\ell_{\mathsf{m},\sigma^{\prime}}(\mathbf{w}\cdot\mathbf{x},y)\right]\leq\mathbb{E}_{(\mathbf{x},y)\sim\mathcal{D}}\left[\ell_{\mathsf{m},\sigma}(\mathbf{w}\cdot\mathbf{x},y)\right]+\frac{\varepsilon}{2}.$

### v2

@@ 4.2 Proof of Theorem 5
$n=\Theta\left(\frac{L^{2}R^{2}}{\varepsilon^{2}}\cdot\left(\frac{\beta^{2}L^{4}R^{4}}{\varepsilon^{2}}\log^{3}\left(\frac{\beta L^{2}R^{2}}{\varepsilon}\right)+\log\left(\frac{1}{\delta}\right)\right)\right),$
@@
$n=\Theta\left(\frac{L^{2}R^{2}}{\varepsilon^{2}}\cdot\left(\frac{\beta L^{2}R^{2}}{\varepsilon}+d\log\left(\frac{\beta L^{2}R^{2}}{\varepsilon}\right)+\log\left(\frac{1}{\delta}\right)\right)\right).$
@@
Throughout this proof, let $\alpha:=\frac{\varepsilon}{6L^{2}R^{2}}$, and let
@@
$n=\Theta\left(\frac{L^{2}R^{2}}{\varepsilon^{2}}\cdot\left(\log\frac{LR}{\delta\varepsilon}+\frac{L^{4}R^{4}\beta^{2}}{\varepsilon^{2}}\log^{3}\left(\frac{L^{3}R^{3}\beta}{\varepsilon^{2}}\right)\right)\right),$
@@
$n=\Theta\left(\frac{L^{2}R^{2}}{\varepsilon^{2}}\cdot\left(\log\frac{LR}{\delta\varepsilon}+\frac{\beta L^{2}R^{2}}{\varepsilon}+d\log\frac{\beta L^{2}R^{2}}{\varepsilon}\right)\right)$
@@ 6.2 Generalization
Let $\mathcal{D}$ be an arbitrary distribution over ×$\mathbb{R}\times\{0,1\}$. For any $\delta\in(0,1/3)$, with probability at least $1-\delta$ over the random draw of $n\geq 2$ i.i.d. data points $(x_{1},y_{1}),\ldots,(x_{n},y_{n})$ from $\mathcal{D}$, the output predictor $p:\mathbb{R}\to[0,1]$ from running PAV on the uniform distribution over the $n$ data points444 When we define PAV in Section 2.4, we assume that the domain $\mathcal{X}$ consists of integers $1,\ldots,n$, but the algorithm extends to any finite domain $\{x_{1},\ldots,x_{n}\}\subseteq\mathbb{R}$ by mapping the elements to $1,\ldots,n$ while preserving the ordering. PAV then gives us a non-decreasing predictor $p:\{x_{1},\ldots,x_{n}\}\to[0,1]$, which we can then extrapolate to a non-decreasing step function (i.e., piece-wise constant with at most $n+1$ pieces) over the entire domain $\mathbb{R}$. satisfies the following properties:
@@ B.1 Omniprediction loss from anti-Lipschitzness
In an instance of Model 1, following the notation (10), let $\sigma\in\mathcal{S}_{0,\beta}$, and let $\alpha\in(0,\frac{1}{2LR})$. There exists a $\sigma^{\prime}\in\mathcal{S}_{\alpha,\alpha+(1-2\alpha LR)\beta}$ such that for all $t\in[-LR,LR]$ and $y\in\{0,1\}$,
$\left|\ell_{\mathsf{m},\sigma^{\prime}}(t,y)-\ell_{\mathsf{m},\sigma}(t,y)\right|\leq\frac{3\alpha L^{2}R^{2}}{2}.$
@@
$\leq\frac{\alpha t^{2}}{2}+\left|\int_{0}^{t}\alpha LR\,\textup{d}\tau\right|\leq\frac{\alpha t^{2}}{2}+\alpha LR\left|t\right|\leq\frac{3L^{2}R^{2}}{2}.$
@@
the conclusion follows from taking expectations over Lemma 24, which gives for our range of $\alpha$,
$\mathbb{E}_{(\mathbf{x},y)\sim\mathcal{D}}\left[\ell_{\mathsf{m},\sigma^{\prime}}(\mathbf{w}\cdot\mathbf{x},y)\right]$ $\leq\mathbb{E}_{(\mathbf{x},y)\sim\mathcal{D}}\left[\ell_{\mathsf{m},\sigma}(\mathbf{w}\cdot\mathbf{x},y)\right]+\frac{\varepsilon}{4},$
$\mathbb{E}_{(\mathbf{x},y)\sim\mathcal{D}}\left[\ell_{\mathsf{m},\sigma^{\prime}}(k_{\sigma^{\prime}}(p(\mathbf{x})),y)\right]$ $\geq\mathbb{E}_{(\mathbf{x},y)\sim\mathcal{D}}\left[\ell_{\mathsf{m},\sigma}(k_{\sigma^{\prime}}(p(\mathbf{x})),y)\right]-\frac{\varepsilon}{4}.\qed$